OpenStudy (anonymous):

Stuck on another problem. Not sure what to do with the formula :/ A trapezoid has an area of 166.5 in.², a height of 9 in., and one base measuring 15 in. What is the length of the other base?

OpenStudy (anonymous):

Tried using the formula but I keep getting stuck

OpenStudy (anonymous):

A= 1/2h(b1+b2)

OpenStudy (anonymous):

Then I came up with 166.5 in^2 = 1/2(9)(15+ b2)

OpenStudy (anonymous):

So 166.5=4.5(15+b2)

OpenStudy (anonymous):

Then what do I do from there :( that's where I'm getting stuck

OpenStudy (anonymous):

I would say distributive property?

OpenStudy (anonymous):

Or divide both sides by 4.5 to give you: 37=15+b2 Subtract 15 from both sides to get b2 by itself: b2=12

OpenStudy (anonymous):

@chrismoon Didn't think of that! That's a good idea.

OpenStudy (anonymous):

Thank you! :) I didn't think of doing it that way cause I thought it wouldn't work

OpenStudy (anonymous):

uh oh lol I just checked the answer choices and I don't see 12 3.5 in. 7 in. 18.5 in. 22 in.

OpenStudy (anonymous):

22 in.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

@chrismoon Or divide both sides by 4.5 to give you: 37=15+b2 Subtract 15 from both sides to get b2 by itself: b2=12 Correct work wrong answer! b2=22 37-15=22!

OpenStudy (anonymous):

Oops, my fault. Must've plugged something into the calculator wrong.